# Ultra-cold atoms as quantum simulators for relativistic phenomena

### v1

@@ 3.3   False-vacuum decay
An important step into the direction of simulating false-vacuum decay in the laboratory is the Trento experiment [261] where an elongated (cigar shaped) spinor Bose-Einstein condensate has been realized. In analogy to the Vienna experiment [222, 223, 224, 225], one may split up the degrees of freedom into the total density $\rho_{+}=\rho_{\uparrow}+\rho_{\downarrow}$ (and phase) components plus density (and phase) differences $\rho_{-}=\rho_{\uparrow}-\rho_{\downarrow}$, which represent the pure spin degrees of freedom. For the latter, we get an effective double-well like potential as in Fig. 3.2 where the energy of the two minima can be changed by varying the detuning of the external microwave radiation which induces Rabi oscillations between the spin states $\uparrow$ and $\downarrow$, similar to the Kibble-Zurek experiment [48] using Ryudberg atoms discussed in the previous section. Due to the strong elongation of the Bose-Einstein condensate, the system is effectively 1+1 dimensional, as in Eq. (3.4). Indeed, bubble formation in analogy to false-vacuum decay has been observed in the Trento experiment [261], but one should be aware of two major deviations from original idea. First, in contrast to the homogeneous case as in Eq. (3.4), the Bose-Einstein condensate is inhomogeneous even in $x$ direction and hence bubbles are formed predominantly at or near the centre of cloud. This is in contrast to spontaneous bubble creation mechanism which can occur anywhere with equal probability [^28] In this regard, the centre of cloud acts a bit like the nucleation seeds (e.g., spoon) for under-cooled water discussed above.. Second, the experimental results and parameters indicate that the false-vacuum decay in the Trento experiment [261] is not caused by quantum fluctuations. Probably it is also not induced by purely thermal (equilibrium) fluctuations, but instead by classical noise generated by the preparation steps. As a result, it is not fully clear (yet) to which extent the quantitative analogy to false-vacuum decay – such as caused by relativistic tunneling in the system (3.4) – can be applied. Nevertheless, the experimental results, such as the observed exponential dependence of the bubble formation probability on experimental parameters over more than one order of magnitude, are a real impressive achievement.
@@
A an alternative realization, one should also mention the Cambridge experiment [262]. In contrast to the scenarios above which are based on weakly interacting condensates, the Cambridge experiment [262] employs strongly interacting (i.e., correlated) bosonic atoms in an optical lattice [^29] Thus, the Cambridge experiment [262] simulating false-vacuum decay displays some similarities to the Sauter-Schwinger or Breit-Wheeler effect in Sec. 2.8 which can also be modeled with strongly interacting fermionic or bosinic atoms in optical lattices, see, e.g., [201, 202] and references therein.. They can be described by an extended Bose-Hubbard Hamiltonian as in Eq. (1.7), but generalized to two atomic species in the presence of an external drive. By resonantly shaking the optical lattice, one can destabilize the Mott insulator state such that it is analogous to the meta-stable false vacuum state, while the true vacuum state (of the driven system) would be a super-fluid state with a staggered order parameter [^30] Such a spatially modulated super-fluid order parameter is often also referred to as super-solid and the associated instability can be visualized via the dispersion relation, where an effective “roton” dip touches the $k$ axis, see, e.g., [263] and references therein.. The meta-stability of the Mott insulator state and its decay to the true ground state have been observed in the Cambridge experiment [262] as a function of the experimental parameters such as the shaking amplitude. Similarly to the Trento experiment [261], the Cambridge experiment [262] is clearly a qualitative analogy to false-vacuum decay (and thus a very important step), but the extent of the quantitative analogy requires further studies.
@@ 3.4   Back-reaction and cosmological constant
A less obvious but very important point where the non-linearity plays a role is the back-reaction of the quantum effects discussed in chapter 2 onto the (approximately classical) background. As an example, let us take the case of Hawking radiation where the in-falling partner particles carry negative energy (from the point of view of a static observer far away). They should decrease the mass of the black hole in order to compensate for the energy emitted via Hawking radiation. However, how this should work in detail is still a matter of debate [^31] This debate is also related to the black-hole information puzzle. On can try to gain some insight using simplified toy models (see, e.g., [269, 270, 271]), but there are many open question..
In this respect, one can hope to learn something from the black-hole analogues, e.g., in Bose-Einstein condensates, because there the microscopic physics is better understood. However, one should also be aware of some important differences between the two cases. First, the fluid analogue such as the de Laval nozzle in Fig. 2.2 is not a closed system – there is a constant in- and out-flux of matter and energy which changes the concept of energy balance in comparison to the real black-hole case. Second, as already discussed above, the black-hole analogues in Bose-Einstein condensates do not reproduce the Einstein equations, which are crucial for the back reaction in real black holes.
Taking into account the above differences, there have been several theoretical investigations regarding the back-reaction for black-hole analogues in Bose-Einstein condensates, see, e.g., [272, 273, 274, 275, 276, 277, 278]. It should also be mentioned here that there were experimental studies [279] of the back-reaction in black-hole analogues – but instead of phonons in Bose-Einstein condensates, surface waves in flowing liquids [158] where used to realize a black-hole analogue (which then typically recovers classical aspects instead of quantum phenomena, at least for the experiments so far).
@@ 4   Summary and outlook
Apart from these experimental aspects, our theoretical understanding of ultra-cold atoms is also quite advanced, e.g., regarding Bose-Einstein condensates in the dilute-gas limit. This combination of experimental and theoretical progress allowed us to propose and even realize several quantum simulators based on ultra-cold atoms – in the spirit of the famous quote from Feynman: “The same equations have the same solutions.” Especially for relativistic phenomena, a selection of proposals, theoretical investigations and experimental realizations have been discussed in the previous sections. In chapter 2, we discussed examples for linear fields – though under the influence of non-trivial external conditions. A recurring theme was the tearing apart of quantum vacuum fluctuations, by gravitational or electric fields, the cosmic expansion or other external influences. Another common feature is quantum squeezing [^35] However, claiming the experimental verification of Hawking radiation or the Unruh effect just because the appearance of squeezing in an experiment is not really justified., which can be visualized by means of the coupled oscillator picture sketched in Fig. 2.1. In chapter 3, we considered some cases of non-linear fields. Here the competition between (local) minima in the potential landscape leads to interesting phenomena beyond standard perturbation theory.
@@ 4.1   Lessons (to be) learned
At the same time, this intuition should be substantiated by hard facts and actual calculations in order to avoid misconceptions and confusion – as Einstein said: “Make things as simple as possible, but no simpler.” In the AI context, training with bad data can lead to wrong recognitions and false classifications. Thus, for each quantum simulator, one must clarify to which extent the analogy applies and where the limits of the analogy are [^38] For example, just finding indications of thermal behaviour and thus concluding (without further arguments) that one found signatures of the Unruh effect or Hawking radiation is probably an exaggeration. As another example, a system where the effective propagation speed vanishes at some position is not necessarily a good analogue for a black hole (where everything can fall in but nothing can escape). As a counter-example, consider the shore line of the ocean where the speed of the water waves goes to zero..
@@ Acknowledgements
I thank A. Weitzel for test reading and help with the figures. Then I would also like to take the opportunity and thank my fellow scientists for many inspiring discussions, where I learned at lot. Even though we often do not agree on everything, an open and constructive scientific debate across all borders (national and other) is extremely valuable. First and foremost, I would like to thank G. Plunien and W. G. Unruh as well as my former and present colleagues from TU Dresden, HZDR, the University of Duisburg-Essen, the University of British Columbia and from the SFB-TR12, SFB1242 and COSLAB. Furthermore, I would like to thank [^39] Most probably, I forgot several important names, for which I apologize – this is no bad intention, but just my poor memory for names. A. Achucarro, I. Affleck, Y. Aharonov, E. Akkermans, A. Andreev, J. Anglin, M. Arndt, M. Aspelmeyer, R. Balbinot, C. Barcelo, I. Bialynicki-Birula, R. Blatt, M. Blencowe, M. Bordag, J. Braden, R. Brandenberger, A. Calogeracos, I. Carusotto, L. Chomaz, M. Choptuik, I. Cirac, E. Cornell, P. Davies, L. Davidovich, B. Dittrich, V. Dodonov, G. Dunne, R. Durrer, F. Dyson, A. Eckardt, J. Edwards, D. Faccio, U. Fischer, S. Floerchinger, L. Ford, K. Fredenhagen, V. Frolov, S. Fulling, L. Garay, B. Geyer, H. Gies, D. Habs, F. Hebenstreit, T. Heinzl, M. Hotta, B.-L. Hu, T. Jacobson, M. Jacquet, M. Johnson, I. Khriplovich, T. Kibble, B. King, A. Kitaev, H. Kleinert, F. Koenig, M. Krusius, T. Kugo, A. Leggett, L. Lehner, U. Leonhardt, S.-P. Kim, S. Liberati, J. Louko, R. Mann, D. Marolf, E. Martín-Martínez, G. Morigi, I. Moss, G. Mourou, V. Mostepanenko, N. Narozhny, M. Oberthaler, D. Page, G. Paraoanu, R. Penrose, T. Pfau, G. Pickett, L. Pitaevskii, S. Popescu, J. Rafelski, M. Raizen, D. Rätzel, J-M. Rost, A. Roura, M. Sakellariadou, R. Sauerbrey, C. Savage, T. Schaetz, J. Schmiedmayer, C. Schubert, M. Scully, G. Semenoff, S. Shankaranarayanan, L. Skrbek, P. Skyba, P. Stamp, J. Steinhauer, G. t’ Hooft, K. Thorne, C. Timm, R. Verch, G. Vidal, M. Visser, G. Vitiello, M. Vojta, G. Volovik, R. Wald, S. Weinfurtner, A. Wipf, R. Woodard, J. Yngvason, A. Zhitnitsky, W. Zurek.

### v2

@@ 3.3   False-vacuum decay
An important step into the direction of simulating false-vacuum decay in the laboratory is the Trento experiment [285] where an elongated (cigar shaped) spinor Bose-Einstein condensate has been realized. In analogy to the Vienna experiment [246, 247, 248, 249], one may split up the degrees of freedom into the total density $\rho_{+}=\rho_{\uparrow}+\rho_{\downarrow}$ (and phase) components plus density (and phase) differences $\rho_{-}=\rho_{\uparrow}-\rho_{\downarrow}$, which represent the pure spin degrees of freedom. For the latter, we get an effective double-well like potential as in Fig. 3.2 where the energy of the two minima can be changed by varying the detuning of the external microwave radiation which induces Rabi oscillations between the spin states $\uparrow$ and $\downarrow$, similar to the Kibble-Zurek experiment [48] using Rydberg atoms discussed in the previous section. Due to the strong elongation of the Bose-Einstein condensate, the system is effectively 1+1 dimensional, as in Eq. (3.4). Indeed, bubble formation in analogy to false-vacuum decay has been observed in the Trento experiment [285], but one should be aware of two major deviations from original idea. First, in contrast to the homogeneous case as in Eq. (3.4), the Bose-Einstein condensate is inhomogeneous even in $x$ direction and hence bubbles are formed predominantly at or near the centre of cloud. This is in contrast to spontaneous bubble creation mechanism which can occur anywhere with equal probability [^38] In this regard, the centre of cloud acts a bit like the nucleation seeds (e.g., spoon) for under-cooled water discussed above.. Second, the experimental results and parameters indicate that the false-vacuum decay in the Trento experiment [285] is not caused by quantum fluctuations. Probably it is also not induced by purely thermal (equilibrium) fluctuations, but instead by classical noise generated by the preparation steps. As a result, it is not fully clear (yet) to which extent the quantitative analogy to false-vacuum decay – such as caused by relativistic tunneling in the system (3.4) – can be applied. Nevertheless, the experimental results, such as the observed exponential dependence of the bubble formation probability on experimental parameters over more than one order of magnitude, are a real impressive achievement.
@@
As an alternative realization, one should also mention the Cambridge experiment [286]. In contrast to the scenarios above which are based on weakly interacting condensates, the Cambridge experiment [286] employs strongly interacting (i.e., correlated) bosonic atoms in an optical lattice [^39] Thus, the Cambridge experiment [286] simulating false-vacuum decay displays some similarities to the Sauter-Schwinger or Breit-Wheeler effect in Sec. 2.8 which can also be modeled with strongly interacting fermionic or bosinic atoms in optical lattices, see, e.g., [223, 224] and references therein.. They can be described by an extended Bose-Hubbard Hamiltonian as in Eq. (1.7), but generalized to two atomic species in the presence of an external drive. By resonantly shaking the optical lattice, one can destabilize the Mott insulator state such that it is analogous to the meta-stable false vacuum state, while the true vacuum state (of the driven system) would be a super-fluid state with a staggered order parameter [^40] Such a spatially modulated super-fluid order parameter is often also referred to as super-solid and the associated instability can be visualized via the dispersion relation, where an effective “roton” dip touches the $k$ axis, see, e.g., [287] and references therein.. The meta-stability of the Mott insulator state and its decay to the true ground state have been observed in the Cambridge experiment [286] as a function of the experimental parameters such as the shaking amplitude. Similarly to the Trento experiment [285], the Cambridge experiment [286] is clearly a qualitative analogy to false-vacuum decay (and thus a very important step), but the extent of the quantitative analogy requires further studies.
@@ 3.4   Back-reaction and cosmological constant
A less obvious but very important point where the non-linearity plays a role is the back-reaction of the quantum effects discussed in chapter 2 onto the (approximately classical) background. As an example, let us take the case of Hawking radiation where the in-falling partner particles carry negative energy (from the point of view of a static observer far away). They should decrease the mass of the black hole in order to compensate for the energy emitted via Hawking radiation. However, how this should work in detail is still a matter of debate [^41] This debate is also related to the black-hole information puzzle. On can try to gain some insight using simplified toy models (see, e.g., [293, 294, 295]), but there are many open question. For the back-reaction in optical systems, see, e.g., [176] and [296]..
In this respect, one can hope to learn something from the black-hole analogues, e.g., in Bose-Einstein condensates, because there the microscopic physics is better understood. However, one should also be aware of some important differences between the two cases. First, the fluid analogue such as the de Laval nozzle in Fig. 2.2 is not a closed system – there is a constant in- and out-flux of matter and energy which changes the concept of energy balance in comparison to the real black-hole case, see also [297]. Second, as already discussed above, the black-hole analogues in Bose-Einstein condensates do not reproduce the Einstein equations, which are crucial for the back reaction in real black holes.
Taking into account the above differences, there have been several theoretical investigations regarding the back-reaction for black-hole analogues and other scenarios in Bose-Einstein condensates, see, e.g., [298, 299, 300, 301, 302, 303, 304, 305, 306]. It should also be mentioned here that there were experimental studies [307] of the back-reaction in black-hole analogues – but instead of phonons in Bose-Einstein condensates, surface waves in flowing liquids [170] where used to realize a black-hole analogue (which then typically recovers classical aspects instead of quantum phenomena, at least for the experiments so far).
@@ 4   Summary and outlook
Apart from these experimental aspects, our theoretical understanding of ultra-cold atoms is also quite advanced, e.g., regarding Bose-Einstein condensates in the dilute-gas limit. This combination of experimental and theoretical progress allowed us to propose and even realize several quantum simulators based on ultra-cold atoms – in the spirit of the famous quote from Feynman: “The same equations have the same solutions” [314]. Especially for relativistic phenomena, a selection of proposals, theoretical investigations and experimental realizations have been discussed in the previous sections. In chapter 2, we discussed examples for linear fields – though under the influence of non-trivial external conditions. A recurring theme was the tearing apart of quantum vacuum fluctuations, by gravitational or electric fields, the cosmic expansion or other external influences. Another common feature is quantum squeezing [^45] However, claiming the experimental verification of Hawking radiation or the Unruh effect just because the appearance of squeezing in an experiment is not really justified., which can be visualized by means of the coupled oscillator picture sketched in Fig. 2.1. In chapter 3, we considered some cases of non-linear fields. Here the competition between (local) minima in the potential landscape leads to interesting phenomena beyond standard perturbation theory.
@@ 4.1   Lessons (to be) learned
At the same time, this intuition should be substantiated by hard facts and actual calculations in order to avoid misconceptions and confusion – as in the quote attributed to Einstein: “Make things as simple as possible, but no simpler.” In the AI context, training with bad data can lead to wrong recognition and false classification. Thus, for each quantum simulator, one must clarify to which extent the analogy applies and where the limits of the analogy are [^48] For example, just finding indications of thermal behaviour and thus concluding (without further arguments) that one found signatures of the Unruh effect or Hawking radiation is probably an exaggeration. As another example, a system where the effective propagation speed vanishes at some position is not necessarily a good analogue for a black hole (where everything can fall in but nothing can escape). As a counter-example, consider the shore line of the ocean where the speed of the water waves goes to zero..
@@ Acknowledgements
I thank A. Weitzel for test reading and help with the figures as well as U. Fischer, T. Jacobson, F. Koenig, S. Liberati, J. Louko, R. Sauerbrey for helpful comments on the manuscript. Then I would also like to take the opportunity and thank my fellow scientists for many inspiring discussions, where I learned at lot. Even though we often do not agree on everything, an open and constructive scientific debate across all borders (national and other) is extremely valuable. First and foremost, I would like to thank G. Plunien and W. G. Unruh as well as my former and present colleagues from TU Dresden, HZDR, the University of Duisburg-Essen, the University of British Columbia and from the SFB-TR12, SFB1242 and COSLAB. Furthermore, I would like to thank [^49] Most probably, I forgot several important names, for which I apologize – this is no bad intention, but just my poor memory for names. A. Achucarro, I. Affleck, Y. Aharonov, E. Akkermans, A. Andreev, J. Anglin, M. Arndt, M. Aspelmeyer, R. Balbinot, C. Barcelo, I. Bialynicki-Birula, R. Blatt, M. Blencowe, M. Bordag, J. Braden, R. Brandenberger, A. Calogeracos, I. Carusotto, L. Chomaz, M. Choptuik, I. Cirac, E. Cornell, P. Davies, L. Davidovich, B. Dittrich, V. Dodonov, G. Dunne, R. Durrer, F. Dyson, A. Eckardt, J. Edwards, D. Faccio, S. Floerchinger, L. Ford, K. Fredenhagen, V. Frolov, S. Fulling, L. Garay, B. Geyer, H. Gies, D. Habs, F. Hebenstreit, T. Heinzl, M. Hotta, B.-L. Hu, M. Jacquet, M. Johnson, I. Khriplovich, T. Kibble, B. King, A. Kitaev, H. Kleinert, M. Krusius, T. Kugo, A. Leggett, L. Lehner, U. Leonhardt, S.-P. Kim, R. Mann, F. Marino, D. Marolf, E. Martín-Martínez, G. Morigi, I. Moss, G. Mourou, V. Mostepanenko, N. Narozhny, M. Oberthaler, D. Page, G. Paraoanu, R. Penrose, T. Pfau, G. Pickett, L. Pitaevskii, S. Popescu, J. Rafelski, M. Raizen, D. Rätzel, J-M. Rost, A. Roura, M. Sakellariadou, C. Savage, T. Schaetz, J. Schmiedmayer, C. Schubert, M. Scully, G. Semenoff, S. Shankaranarayanan, L. Skrbek, P. Skyba, P. Stamp, J. Steinhauer, G. t’ Hooft, K. Thorne, C. Timm, R. Verch, G. Vidal, M. Visser, G. Vitiello, M. Vojta, G. Volovik, R. Wald, S. Weinfurtner, A. Wipf, R. Woodard, J. Yngvason, A. Zhitnitsky, W. Zurek.